# Less is More : A Closer Look at Multi-Modal Few-Shot Learning

### 10 Jan 2024 v1

@@ 4.2 Context-aware Prompt
As have shown in above section, the visual feature can be obtained by a visual backbone for the downstream tasks. However, only considering the visual information is very limited under the few-shot scenarios. To further supplement more available input information, previous works explore additional semantic information as prompts to guide the visual feature backbone. Formally, given a image $\boldsymbol{x}$ from the support set, they pre-define a fixed prompt using the class name of $\boldsymbol{x}$, e.g. cat or a photo of a cat [6]. The pre-defined prompt is related with the class label, denoted as $y^{text}$. Then, the textual semantic feature will be obtained by a pre-trained language model (e.g., CLIP [38]), denoted as $g(\ )$, and formally,
@@
However, these pre-defined fixed prompts may not be flexible enough for various downstream tasks and previous works found that this may lead to the inferior performance [60]. Consequently, we model the input prompts with continuous learnable parameters, which will be optimized in an end-to-end learning way. Specifically, the text prompt $y^{text}$ is represented as a learnable vector $\mathbf{v}$ combined with the class name, which shares with all classes. And the learnable prompt $y^{text}_{v}$ given to the text encoder $g(\ )$ has the following form:
@@
where $v_{i}$ is the $i$-th term of the learnable prompt $\mathbf{v}$, and $L$ denotes the length of context tokens. Specially, the dimension $d_{text}$ of prompt $\mathbf{v}$ keep the same with the language model. In our method, we set $L=4$ and $d_{text}=512$ for all settings and experiments.
@@ 4.3 Multi-modal Feature Fusion
Till now, we have obtained the visual feature $f(x)$ from a visual backbone and the textual semantic feature $g(y^{text}_{\mathbf{v}})$ from a pre-trained language model. Previous works [6, 32] adopt the intricate designed fusion modules to utilize visual and textual information simultaneously. However, these complex fusion methods may ignore the generalization capability of the pre-trained language model and the potential of textual information has been underestimated. Consequently, we just add the visual feature $f(x)$ from a visual backbone and the textual semantic feature $g(y^{text}_{\mathbf{v}})$ directly as the Multi-modal Feature Fusion, and we name this simple baseline as SimpleFSL. Although, there are lots of recent advanced multi-modal fusion mechanisms, we just choose the simplest Add operation in our proposed framework to valide our idea. We argue this simple add operation has little effect with the generalization capability of the pre-trained LM. The more discussions about the alternative multi-modal fusion operations can be found in the experiments.
While there is modal discrepancy between visual and textual representation and the dimensions may be different. Following previous works [6, 32, 51], we adopt an adaptor here to transform the textual representation into the visual representation space and keep the dimension of them to be the same. And we choose a two-layer Multi-Layer Perceptron (MLP) with the non-linear activation function as the adoptor. The details of this adaptor can be found in the section of experiments. This transformation processing can be written as :
$z$ $=adaptor(g(y^{text}_{\mathbf{v}}))$ (4)
$=\boldsymbol{W_{2}}\ \sigma(\boldsymbol{W_{1}}\ g(y^{text}_{\mathbf{v}})+\boldsymbol{b_{1}})+\boldsymbol{b_{2}}$ (5)
@@
Then, we add the visual feature $f(\boldsymbol{x})$ and the transformed textual semantic feature $\boldsymbol{z}$ to compute the prototypes of each classed by averaging the feature of support set within each class [43], and formally we have :
@@
where $\boldsymbol{p_{i}}$ denotes the obtained prototype of $i$-th class. When an unlabeled sample $\boldsymbol{x_{q}}$ from the query set needs to be recognized, we predict it by a non-parametric distance-based classifier with a softmax function:
$\hat{y}_{q}=\frac{\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{i}}\right\rangle/\tau\right)}{\sum_{j}\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{j}}\right\rangle/\tau\right)}$ (7)
@@
$L_{1}=\sum_{D_{base}}CE(\hat{y}_{q},y_{q})$ (8)
where $CE(\ )$ denotes the cross-entropy function.
@@ 4.4 Self-ensemble and Self-Distillation
Besides the introduce SimpleFSL above, we further introduce the Self-ensemble and self-Distillation module based the SimpleFSL to improve the FSL performance further, and we name it as SimpleFSL++. Recall that we have obtained the fusion feature for inference, and we also have obtained the visual feature $f(\boldsymbol{x})$ from the visual backbone, which can also as the input of a classifier, as shown in Figure 2. Similarly, we have:
$\boldsymbol{p_{i}^{0}}$ $=\frac{1}{K}\sum_{j=1}^{K}f(\boldsymbol{x_{j}})$ (9)
$\hat{y}_{q}^{0}=\frac{\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{i}^{0}}\right\rangle/\tau\right)}{\sum_{j}\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{j}^{0}}\right\rangle/\tau\right)}$ (10)
Inspired the research in Knowledge Distillation [18, 58, 57], we ultize the self-ensemble mechanism to build the classifier-3, which ensemble the prediction of $\hat{y}_{q}^{0}$ and $\hat{y}_{q}$, obtained from the classifier-1 and classifier-2 respectively, as shown in Figure 2. In this way, the prediction will be more robust and improved further. Mathematically, the predictions of SimpleFSL++ can be written as:
@@
where $\lambda$ are the weighting factors to balance these two classifer. Additionally, we further utilze the self-Distillation mechanism to let these two classifiers learn reciprocally, which can teach the knowledge and serve as the regularizations mutually [56, 58]. In detail, the self-distillation mechanism employed can be described as:
@@
where KL denotes the the knowledge distillation loss, and we opt Kullback-Leibler Divergence as our implement.
In summary, the final loss of the proposed SimpleFSL++ can be described as:
@@
where the $L_{2}=\sum_{D_{base}}CE(\hat{y}_{q}^{0},y_{q})$ denotes the cross-entropy loss for the classifier-2, and $\alpha$ is a hyper-parameters.
@@ 5 Experiments
In this section, we evaluate our proposed SimpleFSL and SimpleFSL++ framework on widely used few-shot learning benchmarks and compare them with current state-of-the-art methods.
@@
We conduct the FSL experiments on four image datasets: miniImageNet [47], tieredImageNet [39], CIFAR-FS [9, 24] and FC100 [35]. The first two are all the subset of ILSVRC-12 dataset [40], and the last two are all derived from the CIFAR-100 dataset [24]. These datasets are all public available. MiniImageNet: It contains 100 classes and 60,000 images, in which 64 classes are assigned to the base dataset, 16 classes to the validation dataset, and 20 classes to the novel dataset, respectively. TieredImageNet: It is a larger dataset compared to the MiniImageNet, containing 608 classes and 779,165 images, and the base and novel datasets of it is more semantically different. We follow the previous split proposed by [39], in which 351, 97, and 160 classes are used for the base, validation, and novel datasets, respectively. CIFAR-FS: It containing 100 classes and 600 images per class. Following previous works [9], we use 64 classes to construct the base dataset, 16 classes for the validation dataset, and the remaining 20 classes for the novel dataset. FC100: It also contains 60,000 images , with 100 classes, and was split into 60 training classes, 20 validation classes and 20 test classes, according their semantic superclasses. Consequently, the large semantic gap makes FC100 more difficult than CIFAR-FS.
@@
SimpleFSL++ (Ours) Visformer-T 10.0M/$1.3\times 10^{9}$ $\mathbf{75.27}\pm\mathbf{0.71}$ $\mathbf{83.89}\pm\mathbf{0.54}$ $\mathbf{80.52}\pm\mathbf{0.81}$ $\mathbf{88.36}\pm\mathbf{0.60}$
[…]
Implementation details. Our proposed SimpleFSL and Simple++ both contain the visual backbone and textual backbone for the feature extraction. For the visual branch, we opt the Visformer-Tiny [8] as the default visual backbone, and resize the all input image with 224×224 pixel. Compared to the usually used ResNet-12 [16], the Visformer-Tiny has the similar number of parameters but with less floating point operations (FLOPS). And Unless otherwise specified, the dimension of obtained visual representation is 384. For the textual branch, we adopt the text encoder of pre-trained CLIP [38] as the default textual backbone, due to the fact that it is available for the learnable prompt tuning, as described above. Specially, we do not use the visual encoder of CLIP for a fair comparison, and its code and pre-trained weights are available for public use  [^2] https://github.com/openai/CLIP.. Specially, we freeze the all parameters of textual backbone and only train other parameters in our two frameworks. we use the embeddings of “a photo of a” to initialize the learnable prompt $\mathbf{v}$, and the dimension of obtained textual representation is all 512.
In pre-training stage and meta-traing stage, we all employ the AdamW optimer [33] with a learning rate of 5e-4 and a weight decay of 5e-2. Specially, we reduce the learning rate of the visual backbone to 1e-6 during the meta-training stage, while others keep ohers the same. For evaluation, we test our method in 5 way-1 shot/5 shot settings on the novel dataset and randomly sample 2,000 few-shot tasks from it. Then, we report the top-1 mean accuracy (%) with the 95% confidence interval. It worth to note that our proposed frameworks both do not need to fine-tune during the evaluation stage. All experients are conducted on a Linux machine with a single NVIDIA RTX3090 GPU.
@@ 5.2 Main results
SimpleFSL++ (Ours) Visformer-T 10.0M/$1.3\times 10^{9}$ $\mathbf{84.96}\pm\mathbf{0.64}$ $\mathbf{89.10}\pm\mathbf{0.31}$ $\mathbf{49.37}\pm\mathbf{0.64}$ $\mathbf{60.07}\pm\mathbf{0.65}$
Table 1 and Table 2 summarize the performances of our proposed framework SimpleFSL and SimpleFSL++ with recent state-of-the-art FSL methods under the 5-way 1/5-shot learning tasks under four datasets. The compared FSL methods include both the single-modal based FSL (only use the visual information) and multi-modal based FSL (use the textual information of class). Firstly, observed from these these experimental results in two tables, our simple simpleFSL and simpleFSL++ achieve satisfactory performances in all four FSL datasets. Especially in the 5-way 1-shot setting, our SimpleFSL and SimpleFSL++ both surpass the SP-CLIP [6] with the obvious improvements. For example, SimpleFSL++ achieve the 3.9% relative accuracy improvement compared the SP-CLIP on miniImageNet. SimpleFSL achieve the 3.2% accuracy improvement compared the SP-CLIP on CIFAR. We argue these obvious performance improvements should be attributed to that we explicitly utilize the generalization capacity of the pre-trained LM coordinating with the more flexible learnable prompts, thought the frameworks of our proposed SimpleFSL and SimpleFSL++ are both simple. Secondly, the proposed SimpleFSL++ can beat the SimpleFSL achieve better classification accuracy on four datasets. The details analysis can be found in Ablation study. Thirdly, we observe that our SimpleFSL and SimpleFSL++ both have more preponderant in 1-shot classification tasks than 5-shot classification tasks. We argue that the visual supervision signal in 5-shot learning is more sufficient than 1-shot learning , and dominate the model training [51, 6].
@@ 5.3.1 Ablation study
+ fixed Prompt 72.78 80.21 81.12 46.46
+ Learnable Prompt 74.80 80.06 84.81 48.77
@@
+ self-Distillation 75.27 80.52 84.96 49.37
@@
Dateset-aware $\mathbf{75.27}$ $\mathbf{83.89}$ $\mathbf{84.92}$ $\mathbf{89.07}$

### 24 Mar 2024 v2

@@ 4.2 Context-aware Prompt
As illustrated in the preceding section, the visual feature can be obtained by a visual backbone for the downstream tasks. Notwithstanding, reliance solely on the visual information may cause insufficient supervision signals and suboptimal performance [58]. To supplement more available input information, previous works explore additional semantic information as the prompts to guide the visual feature extraction. Formally, given an image $\boldsymbol{x}$ from the support set, they pre-define a fixed prompt using the class name of $\boldsymbol{x}$, e.g. cat [58, 34] or a photo of a cat [6]. The pre-defined prompt, denoted as $y^{text}$, is related with the class label. Then, the textual semantic feature will be procured by an off-the-shelf pre-trained language model (e.g., CLIP [44]), denoted as $g(\ )$, formulated as:
@@
However, these pre-defined fixed prompts may not be flexible enough for various downstream tasks and previous works found that this may lead to the inferior performance [68]. Consequently, we address this by modeling the input prompts with continuous learnable parameters, which will be optimized in the end-to-end learning way, thereby allowing for flexible accommodation to various datasets. Specifically, the text prompt $y^{text}$ is instantiated as a learnable vector $\mathbf{v}$ combined with the class name and shared with all classes in our implements. The learnable prompt $y^{text}_{v}$ fed into the text encoder $g(\ )$ takes the form:
@@
where $v_{i}$ is the $i$-th term of the learnable prompt $\mathbf{v}$, and $L$ denotes the length of context tokens. Specially, the dimension $d_{text}$ of prompt $\mathbf{v}$ keep the same with the language model. In our method, we set $L=4$ and $d_{text}=512$ across all experimental setups.
@@ 4.3 Multi-modal Feature Fusion
Thus far, we have extracted the visual feature $f(x)$ from a visual backbone and the textual semantic feature $g(y^{text}_{\mathbf{v}})$ from a pre-trained LM. Previous works [6, 34] have employed the intricate designed fusion modules to utilize visual and textual information simultaneously. However, these complex fusion methods may inadvertently compromise the generalization capability of the pre-trained LM. Consequently , the potential of semantic information for FSL has been underestimated. To address this, we straightforwardly add the visual feature $f(x)$ and the textual semantic feature $g(y^{text}_{\mathbf{v}})$ as the Multi-modal Feature Fusion, and we name this simple baseline as SimpleFSL. Despite the availability of advanced multi-modal fusion techniques [15], we choose the simplest Add operation in our proposed framework to demonstrate our concept. We argue this simple add operation has little effect with the generalization capability of the pre-trained LM. The more discussions about the alternative multi-modal fusion operations can be found in the experiments section.
While there is the inherent discrepancy between visual and textual modalities, including differing dimensions. Following previous works [6, 34, 58], we adopt an adaptor here to transform the semantic feature space into the visual feature space, and ensure dimensional consistency. We choose a simple two-layer Multi-Layer Perceptron (MLP) with the non-linear activation function as the adaptor in our implements. The discussion about other adaptors can be found in the experiments section. The transformation process is expressed as:
$\boldsymbol{z}$ $=adaptor(g(y^{text}_{\mathbf{v}}))=\boldsymbol{W_{2}}\ \sigma(\boldsymbol{W_{1}}\ g(y^{text}_{\mathbf{v}})+\boldsymbol{b_{1}})+\boldsymbol{b_{2}}$ (4)
@@
Subsequently, we compute the class prototypes [49] by averaging the sum of visual features $f(\boldsymbol{x})$ and the transformed semantic feature $\boldsymbol{z}$ for each sample in the support set, formulated as:
@@
where $\boldsymbol{p_{i}}$ denotes the obtained prototype of $i$-th class. Classification of an unlabeled sample $\boldsymbol{x_{q}}$ from the query set is performed using a non-parametric distance-based classifier [7, 49] with a softmax function:
$\hat{y}_{q}=\frac{\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{i}}\right\rangle\right)}{\sum_{j}\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{j}}\right\rangle\right)}$ (6)
@@ 4.4 Self-ensemble and Self-Distillation
Building upon SimpleFSL, we introduce SimpleFSL++, which incorporates self-ensemble and self-distillation modules to further refine FSL performance. Recall that we have obtained the fusion feature for inference, and we also have obtained the visual feature $f(\boldsymbol{x})$ from the visual backbone, which can also serve as the input of a classifier, as shown in Fig.2. Similarly, we have: $\boldsymbol{p_{i}^{0}}=\frac{1}{K}\sum_{j=1}^{K}f(\boldsymbol{x_{j}})$ and $\hat{y}_{q}^{0}=\frac{\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{i}^{0}}\right\rangle/\tau\right)}{\sum_{j}\exp\left(\left\langle f(\boldsymbol{x_{q}}),\boldsymbol{p_{j}^{0}}\right\rangle/\tau\right)}$.
Inspired the research in Knowledge Distillation [19, 66, 65], we utilize the self-ensemble mechanism to build the classifier-3, which ensembles the prediction of $\hat{y}_{q}^{0}$ and $\hat{y}_{q}$, obtained from classifier-1 and classifier-2 respectively, as shown in Figure 2. In this way, the prediction will be more robust and improved further. Mathematically, the predictions of SimpleFSL++ is written as:
@@
where $\lambda$ is the weighting factor to balance these two classifiers. Additionally, we further utilize the self-Distillation mechanism to allow these classifiers learn reciprocally, which can transfer the learned knowledge and serve as the regularization mutually [64, 66]. In detail, the self-distillation mechanism employed is formulated as:
@@
where $KL$ refers to the knowledge distillation loss, and we opt Kullback-Leibler Divergence as our implementation.
To summarize, the comprehensive loss of the proposed SimpleFSL++ is articulated as:
@@
where $L_{2}=\sum_{D_{base}}CE(\hat{y}_{q}^{0},y_{q})$ represents the cross-entropy loss for the classifier-2, and $\alpha$ is a hyper-parameters. Additionally, it is worth to note that self-distillation is not required during inference.
@@ 5 Experiments
We conduct the FSL experiments on four widely used datasets: miniImageNet [54], tieredImageNet [45], CIFAR-FS [9, 25] and FC100 [39]. The first two are subsets of ILSVRC-12 dataset [46], and the last two derive from the CIFAR-100 dataset [25]. These datasets are all public available. MiniImageNet: This dataset comprises 100 classes and 60,000 images, in which 64 classes are assigned to the base dataset, 16 classes to the validation dataset, and 20 classes to the novel dataset. TieredImageNet: It is a larger dataset compared to MiniImageNet, containing 608 classes and 779,165 images, and the base and novel datasets of it is more semantically different. We follow the previous split proposed by [45], in which 351, 97 and 160 classes are used for the base, validation, and novel datasets, respectively. CIFAR-FS: It contains 100 classes and 600 images per class. Following previous works [9], we use 64 classes for the base dataset, 16 classes for the validation dataset, and the remaining 20 classes for the novel dataset. FC100: It also contains 60,000 images , with 100 classes, and was split into 60 base classes, 20 validation classes and 20 novel classes, according their semantic superclasses. Consequently, its discernible semantic gap presents a steeper challenge compared to CIFAR-FS.
[…]
@@
Meta-AdaM [51] ResNet-12 12.5M/$3.5\times 10^{9}$ $59.89\pm 0.49$ $77.92\pm 0.43$ $65.31\pm 0.48$ $85.24\pm 0.35$
CORL [16] ResNet-12 12.5M/$3.5\times 10^{9}$ $65.74\pm 0.53$ $83.03\pm 0.33$ $73.82\pm 0.58$ $86.76\pm 0.53$
@@
SimpleFSL++ (Ours) Visformer-T 10.0M/$1.3\times 10^{9}$ $\mathbf{75.59}\pm\mathbf{0.37}$ $\mathbf{83.89}\pm\mathbf{0.54}$ $\mathbf{80.52}\pm\mathbf{0.81}$ $\mathbf{88.36}\pm\mathbf{0.60}$
Our proposed SimpleFSL and SimpleFSL++ both contain the visual backbone and textual backbone for the feature extraction. we opt the Visformer-Tiny [8] as the visual backbone, and resize the all input image with 224×224 pixel. Compared to the usually used ResNet-12 [18], the Visformer-Tiny has the similar number of parameters but with less floating point operations (FLOPS). And unless otherwise specified, the dimension of obtained visual representation is 384. we adopt the text encoder of pre-trained CLIP [44] as the textual backbone, due to the fact that it has distinguish performance [6] and supports the learnable prompt tuning, as described above. The code and pre-trained weights of CLIP are available for public use  [^2] https://github.com/openai/CLIP.. Notably, we do not use the visual encoder of CLIP for a fair comparison. During the training stage, we freeze the all parameters of textual backbone and optimize only the other model parameters. We adopt the embeddings of “a photo of a” to initialize the learnable prompt $\mathbf{v}$, and the dimension of obtained textual representation is consistently 512.
In pre-training and meta-training stages, we all employ the AdamW optimer [36] with a learning rate of 5e-4 and a weight decay of 5e-2. Specially, we reduce the learning rate of the visual backbone to 1e-6 during the meta-training stage, keeping others the same. For evaluation, we test our framework under 5 way-1 shot/5 shot settings on the novel dataset and randomly sample 2,000 few-shot tasks from it. Then, we report the top-1 mean accuracy (%) with the 95% confidence interval. Notably, our proposed frameworks both do not require fine-tuning during evaluation, while some FSL baselines require [2, 9, 13]. All experiments are conducted on a Linux machine with a single NVIDIA RTX3090 GPU.
@@ 5.2 Main results
Meta-AdaM [51] ResNet-12 12.5M/$3.5\times 10^{9}$ $41.12\pm 0.49$ $56.14\pm 0.49$
LEP-CLIP [60] ResNet-12 12.5M/$3.5\times 10^{9}$ $80.62\pm 0.41$ $86.22\pm 0.33$
@@
SimpleFSL++ (Ours) Visformer-T 10.0M/$1.3\times 10^{9}$ $\mathbf{85.09}\pm\mathbf{0.64}$ $\mathbf{89.10}\pm\mathbf{0.31}$ $\mathbf{49.37}\pm\mathbf{0.64}$ $\mathbf{60.07}\pm\mathbf{0.65}$
Table 1 and Table 2 summarize the performances of our proposed two framework SimpleFSL and SimpleFSL++ compared to recent state-of-the-art FSL methods under 5-way 1/5-shot learning tasks on the aforementioned four datasets. The compared FSL methods include both the single-modal based FSL (visual information only) and multi-modal based FSL (including semantic information of classes). Firstly, observed from these experimental results in two tables, our simpleFSL and simpleFSL++ achieve satisfactory performances on all FSL datasets. Especially in the 5-way 1-shot setting, our SimpleFSL and SimpleFSL++ both surpass the SOTA SP-CLIP [6] and LEP-CLIP [60] with substantial accuracy gains. For example, SimpleFSL++ achieve a 4.4% relative accuracy improvement over the SP-CLIP on miniImageNet, and SimpleFSL achieve a 3.2% relative accuracy improvement on CIFAR. These obvious performance improvements should be attributed to the explicit utilization of the pre-trained LM’s generalization capacity coordinating with the adaptable learnable prompts, despite the simplicity of both SimpleFSL and SimpleFSL++ frameworks without complex and sophisticated structures. In contrast, previous semantic-based baselines predominantly focus on designing the sophisticated fusion mechanism, ignoring the explicit utilization of LMs. Secondly, the proposed SimpleFSL++ consistently outperforms SimpleFSL with better classification accuracy on four datasets, derived from its self-ensemble and self-distillation mechanisms. The detail analysis can be found in Ablation study. Thirdly, our evaluation indicates that our SimpleFSL and SimpleFSL++ both have a more pronounced edge in 1-shot than 5-shot classification tasks. This suggests that the visual supervision signal in 5-shot learning is more abundant than 1-shot learning , and dominate the model training compared to the semantic supervision signal. These observations align with insights from previous research [58, 6].
@@ 5.3.1 Ablation study
+ fixed Prompt 72.78 80.06 81.12 46.46
+ Learnable Prompt 74.80 80.21 84.81 48.77
@@
+ self-Distillation 75.59 80.52 85.09 49.37
Fig.3 presents the results of the ablation study conducted on four datasets under the 5-way 1-shot learning setting. By combining the visual backbone with the pre-trained LM even with the fixed Prompts, the accuracy can be improved notably. Subsequently, employing the learnable prompts instead of fixed prompts further amplifies performance gains with considerable margins. Moreover, the incorporation of self-ensemble and self-distillation module in SimpleFSL++ further elevates the accuracy on all four datasets. Collectively, the proposed modules used in SimpleFSL and SimpleFSL++ are instrumental to performance. The ablation study underscores the significance of explicitly leveraging pre-trained LMs for FSL, which can lead to substantial improvements.
5.3.2 Prompts analysis
@@
Dateset-aware $\mathbf{75.59}$ $\mathbf{83.89}$ $\mathbf{85.09}$ $\mathbf{89.10}$
@@
Building on the aforementioned discussion, we introduce the learnable prompts instead of previous fiexed prompts in our framework, and ablation study demonstrates its effectiveness for FSL. The learnable prompts we used in SimpleFSL and SimpleFSL++ are all dataset-aware, meaning the learned vector are shared with all classes in a dataset. Inspired by recent progresses in prompt Learning [67], we evaluate the class-aware prompts, which are related to specific classes and conditional on the visual prototypes. Further, considering the fact of that few-shot learning consists of many few-shot tasks, we propose task-aware prompts, designed to be unique to each few-shot task and conditional on the visual features of all samples in a FSL task. The more experimental details and settings about the designs of class-aware prompt and task-aware prompt can be found in Appendix. Fig.4 summarize the 5-way 1/5-shot accuracy comparisons of the three learnable prompt variants. And in Fig.4, the fine granularity of modeling decreases from top to bottom sequentially. Unfortunately, the deployments of task-aware or class-aware prompts both do not lead to better performance. These phenomena consist with the conclusion in previous research [68], which suggests that more fine-grained modeling of prompts might not necessarily enhance downstream task performance. And these potentially are caused by overfitting because the more fine-grained modeling, the more parameters to learning, which is particularly challenging in low-data scenarios. Consequently, we adopt the dataset-aware prompts in our framework and the quest for designing more effective prompts for FSL is a promising avenue for future inquiry.
5.3.3 Adaptor analysis